Find the mean and variance for the following probability distribution
X -1 -2 3 4
P(X) 0.05 0.43 0.32 0.20

Answer by rothauserc(4718)

mean of x = expected value of x = summation i = 1 to 4 of x(i) * p(x(i))
:
mean = -1(0.05) + -2(0.43) + 3(0.32) + 4(0.20) = 0.85
:
variance = summation i = 1 to 4 of (x(i) - mean)^2) * p(x(i))
:
variance = ((-1 - 0.85)^2 * (0.05)) + ((-2 - 0.85)^2 * (0.43)) + ((3 - 0.85)^2 * (0.32)) + ((4 - 0.85)^2 * (0.20)) = 7.1275
: